OpenStudy (anonymous):

does anyone take Algebra2 in connections that's could help me? im stuck :c

OpenStudy (anonymous):

Yeah, I do. What you need?

OpenStudy (anonymous):

I need help with a few question on Unit 2 Lesson 10

OpenStudy (anonymous):

I'm in unit 3 lesson 5 so maybe I could help o. o

OpenStudy (anonymous):

Ok. What exactly in the test do you need help with?

OpenStudy (anonymous):

2. What is (are) the x-intercept(s) of the function graphed above? a.-4 b.-14 ad 4.5 c.-5, 1, and 4 d.-4, -1, and 5

OpenStudy (anonymous):

D. Points -4 -1 and 5 are all on the x-graph.

OpenStudy (anonymous):

Just realized the test is randomized, so none of these were in my test lol

OpenStudy (anonymous):

thank you :)! and really i thought they were all the same lol

OpenStudy (anonymous):

Np :) anything else?

OpenStudy (anonymous):

i do have a few others if you don't mind s:

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

What is f(-3) if f(x)= l 2x+1l + 5? a.0 b.10 c. 12 d.-10

OpenStudy (anonymous):

C, 12. Plug in -3 for x :)

OpenStudy (anonymous):

Wait. I had a similar problem and got it wrong. It's b, 10. Trick question lol

OpenStudy (anonymous):

ohhh! okay lol

OpenStudy (anonymous):

do you mind helping me with 1 more sorry .-.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x: 1, 2, 3, 4, 5, 6, 7 y: 13,17,14,8,3,-2,2 Which equation represents the regression line for the data given in the table above? a.y=-3x-20 b.y=-3x+20 c.y=3x+20 d.y=3x-20

OpenStudy (anonymous):

B, its going down so its negative 3. It starts at y-intercept 20 so its +20.

OpenStudy (anonymous):

desmos.com is a good place to graph the plots btw :)

OpenStudy (anonymous):

thank you :) im gonna try it out!

OpenStudy (anonymous):

Yw :) I gotta go; good luck with your test :)

OpenStudy (anonymous):

it does work really well ^_^ thanks! i will make another post lol I'm on my last question ill tag you in it and give you a medal c: ty for helping me so much i appreciate it